# Invariant sets for homeomorphisms of hyperbolic 3-manifolds

### v1

@@ Abstract.
Keywords: Quasigeodesic flows, 3-manifolds, foliations, rotation theory, minimal homeomorphisms. MSC 2022: Primary: 57R30, 37C27; Secondary: 53C12, 53C23.
@@ 1. Introduction
Let $f:M\to M$ be a minimal homeomorphism of a closed hyperbolic 3-manifold homotopic to the identity. Then, there is a $G_{\delta}$-dense111i.e. containing a countable intersection of open and dense sets. set of points $\mathcal{G}\subset M$ so that if $x\in\mathcal{G}$ and $\tilde{x}\in\widetilde{M}$ is a lift of $x$ then:
@@
For flows, in [Fra] more information than that given by Theorem A is obtained about these compact sets (one can show these are periodic orbits of the flow), but such a result cannot hold for homeomorphisms (see Example 5.27). One can still expect to say more about the structure of these sets, but we achieve this only for some particular class of quasi-geodesic homeomorphisms. The following result is the technical core of this paper and extends ideas going back to [Ha] which were developed in some particular cases in [BFFP].
@@
Here we need to precise some terms. Given a foliation $\mathcal{F}$ of a closed 3-manifold, we say it is $-\mathbb{R}$-covered if when lifted to the universal cover, we have that $\widetilde{\mathcal{F}}$ verifies that every leaf is a properly embedded plane in $\widetilde{M}$ and for every pair of leaves $L,L^{\prime}$ we have that the Hausdorff distance between $L$ and $L^{\prime}$ is bounded (see [FP1] for discussion, and below we give some more equivalences of this notion). Such foliations are quite abundant in hyperbolic 3-manifolds and extend the notion of fibrations for fibered hyperbolic 3-manifolds: they are always blow ups of slitherings (see [Thu2, Cal2]). The reason for calling them $\mathbb{R}$-covered is because in particular, we can see that the leaf space $\mathcal{L}_{\mathcal{F}}=\widetilde{M}/_{\mathcal{F}}$ is homeomorphic to $\mathbb{R}$.
We say that $f$ has positive escape rate with respect to $\mathcal{F}$ if for every $x\in\widetilde{M}$ and a good lift $\widetilde{f}$ of $f$ we have that the leaf of $\widetilde{f}^{n}(x)$ goes to +$\infty$+ ∞ in $\mathcal{L}_{\mathcal{F}}$.
As an application of Theorem C, in § 6 we show that if a homeomorphism has positive speed with respect to some homological direction then it must have many invariant sets.
@@
Acknowledgements: Part of this paper is the content of the master thesis [Go] of the first author, made in PEDECIBA-Udelar in Uruguay. The authors would like to thank Ian Agol, Alfonso Artigue, Jairo Bochi, Sylvain Crovisier, Sergio Fenley, Pablo Lessa, Ana Rechtman, Jonathan Zung for useful comments and exchange.
@@ 2.1. Escape rate
[…]
@@ 2.2. Homological rotation
The proof of this theorem is an application of Theorem C together with some properties of surfaces in hyperbolic 3-manifolds.
@@
Let $f:M\to M$ be a uniquely ergodic homeomorphism of a closed hyperbolic manifold homotopic to the identity, then, for every $\pi:\hat{M}\to M$ finite cover, we have that the lift $\hat{f}:\hat{M}\to\hat{M}$ satisfies that for every $c\in H^{1}(\hat{M},\mathbb{R})$, if $\hat{\mu}$ is the lift of the unique invariant measure of $\mu$, then $\ell_{c}(\mu)=0$.
@@
Let $f:M\to M$ be a homeomorphism which escapes to infinity with uniform positive rate [^2] I.e. there exists $\tau>0$ such that $\ell_{\mu}\geq\tau$ for every $f$-invariant ergodic measure $\mu$.. Is it true that there exists a finite cover $\hat{M}$ of $M$ such that for the lift $\hat{f}$ of $f$ to $\hat{M}$ there is a cohomology class $c\in H^{1}(\hat{M},\mathbb{R})$ for which $\hat{f}$ has uniform positive escape rate [^3] I.e. for every $\hat{f}$-invariant ergodic measure $\hat{\mu}$ one has that $\ell_{c}(\hat{\mu})>0$. with respect to $c$?
@@
If $f:M\to M$ is a minimal homeomorphism of a closed $3$-manifold homotopic to the identity and $\widetilde{f}:\widetilde{M}\to\widetilde{M}$ is a good lift, then, for every $x\in\widetilde{M}$ we have that the sequence $d(\widetilde{f}^{n}(x),x)$ is unbounded. In fact, if for some point $x$ this sequence is bounded, we can consider the closure of its $\widetilde{f}$-orbit, the boundary of which is a compact $\widetilde{f}$-invariant set with empty interior. Thus, its projection to $M$ is a closed, proper, $f$-invariant set, contradicting minimality.
@@ 2.3. The quasi-geodesic case
This will be proved in § 3.
@@ 2.4. $\mathbb{R}$-covered foliations
$B(Y,\varepsilon)$ :=$:=\{x\in\widetilde{M}:d(x,Y)<\varepsilon\}$
$B_{L}(X,\varepsilon)$ :=$:=\{x\in L:d_{L}(x,X)<\varepsilon\}.$
If the foliation has no torus leaves (more generally, if it is Reebless) we know that the leaf space $\mathcal{L}=\widetilde{M}/_{\widetilde{\mathcal{F}}}$ is a simply connected (possibly non-Hausdorff) one-dimensional manifold. We say that $\mathcal{F}$ is $\mathbb{R}$-covered if $\mathcal{L}$ is Hausdorff (equivalently, homeomorphic to $\mathbb{R}$). We say that $\mathcal{F}$ is uniform if for every pair of leaves $L,L^{\prime}\in\widetilde{\mathcal{F}}$ there is $C>0$ such that $L$ is contained in the $C$-neighborhood of $L^{\prime}$ and viceversa. See [Cal2, Chapter 9] for more on these foliations, in particular we will use [Cal2, Thm. 9.15 and Lemma 9.10].
@@
If $\widetilde{\mathcal{F}}$ is $\mathbb{R}$-covered and uniform, there exists a homeomorphism $Z:\mathcal{L}\to\mathcal{L}$ that commutes with the action of $\pi_{1}(M)$ on $\mathcal{L}=\widetilde{M}/\widetilde{\mathcal{F}}$, and a constant $c>0$ such that for every leaf $L\in\widetilde{\mathcal{F}}$, we have the bound $d(L,Z(L))>c$.
@@ 3. Quasi-geodesic homeomorphisms
In this section we consider a homeomorphism $f:M\to M$ of a hyperbolic 3-manifold which is homotopic to the identity and such that it is quasi-geodesic in the sense of §2.3. The goal of this section is to prove Theorem 2.8 and that such homeomorphisms must have positive topological entropy. Together, this will imply Theorem A.
@@ 3.1. Constructing the invariant set
Note that we do not claim that the image of these maps is closed, so they need not be surjective. These properties are analogous to some properties verified for quasi-geodesic flows shown in [Cal2, §10.7].
@@
This is a direct consequence of the properties we proved for the functions $e^{+}$ and $e^{-}$ in Lemmas 3.1 and 3.3. ∎
@@ 3.2. Invariant subsets
Since $\gamma_{n}\to\infty$ in $\pi_{1}(M)$ (this is equivalent to saying that the norm $\gamma_{n}\in\mathrm{Isom}(\mathbb{H}^{3})\cong\mathrm{PSL}_{2}(\mathbb{C})$ goes to infinity) we know that there is a sequence of neighborhoods $U_{n}\subset\partial\mathbb{H}^{3}$ shrinking to $v_{+}$ (with the induced metric of $\mathbb{R}^{3}$) whose complement is mapped in some open set $V_{n}$ of diameter going to 00 (again, with the induced metric of $\mathbb{R}^{3}$). We can choose also $U_{n}$ and $V_{n}$ so that the action of $\gamma_{n}$ expands all vectors tangent to $U_{n}$. Also, since $v_{+}\neq v_{-}$, for large enough $n$ we know that $v_{-}\notin U_{n}$. By cutting $\eta_{0}$ if necessary, we can assume that $e^{-}(\eta_{0}([0,1])\cap U_{n}=\emptyset$. Up to taking a subsequence, we know that $V_{n}\to\xi\in\partial\mathbb{H}^{3}$ and thus we get that $\gamma_{n}e^{-}(\eta_{0}([0,1]))=e^{-}(\gamma_{n}\eta_{0}([0,1]))\to\xi$.
@@
Let $D_{i}:=\mathcal{W}^{u}_{10\epsilon}(\mathcal{W}^{s}_{10\epsilon}(x_{i}))$ for every $i\in\{1,\ldots,N\}$, for some $\epsilon\in(0,\epsilon_{0})$ small enough so that if $i\neq j$, then $D_{i}\cap D_{j}=\emptyset$ and, moreover, if a point $x\in D_{i}$ satisfies that $G_{t}(x)\in D_{j}$ for some $t>0$, then $t\geq T$. Moreover, let $\epsilon\in(0,\epsilon_{0})$ be small enough so that $I$ admits a sub-chainable continua $J_{1}$ of diameter $2\epsilon$. Also let $D^{\prime}_{i}:=\mathcal{W}^{u}_{\epsilon/2}(\mathcal{W}^{s}_{\epsilon/2}(x_{i}))$ for every $i\in\{1,\ldots,N\}$.
@@
Note that $\operatorname{diam}(G_{t}(J_{1}))$ tends to infinity with $t$ since $J_{1}$ is contained in a strong unstable manifold of $G_{t}$. Let $t_{1}>0$ be the first positive time such that one of the following two conditions happen: either $\operatorname{diam}(G_{t_{1}}(J_{1}))=3\epsilon$ or $G_{t_{1}}(J_{1})\cap D^{\prime}_{i_{1}}\neq\emptyset$ for some ${i_{1}}\in\{1,\ldots,N\}$. If the former happens, let $J_{2}\subset G_{t_{1}}(J_{1})$ be a sub-chainable continua such that $\operatorname{diam}(J_{2})=2\epsilon$. If the latter happens, note that $G_{t_{1}}(J_{1})\subset D_{i_{1}}$, and let $J_{2}\subset G_{t_{1}}(J_{1})$ be a sub-chainable continua disjoint from $D_{i_{1}}^{\prime}$ such that $\operatorname{diam}(J_{2})=\epsilon$. Let $I_{1}=J_{1}$ and $I_{2}:=G_{-t_{1}}(J_{2})$. Note that $I_{2}\subset I_{1}$.
@@
Now, let $t_{2}>0$ be the first positive time such that, again, one of the following two conditions happen: either $\operatorname{diam}(G_{t_{2}}(J_{2}))=3\epsilon$ or $G_{t_{2}}(J_{2})\cap D^{\prime}_{i_{2}}\neq\emptyset$ for some ${i_{2}}\in\{1,\ldots,N\}$. If the former happens, let $J_{3}\subset G_{t_{2}}(J_{2})$ be a sub-chainable continua such that $\operatorname{diam}(J_{3})=2\epsilon$. If the latter happens, note that $G_{t_{2}}(J_{2})\subset D_{i_{2}}$, and let $J_{3}\subset G_{t_{2}}(J_{2})$ be a sub-chainable continua disjoint from $D^{\prime}_{i_{2}}$ such that $\operatorname{diam}(J_{2})=\epsilon$ (note that in this case, this is possible since $\operatorname{diam}(G_{t_{2}}(J_{2}))\geq 2\epsilon$ is guaranteeed because of Claim 3.10). Let $I_{3}:=G_{-t_{1}-t_{2}}(J_{3})$. Note that $I_{3}\subset I_{2}\subset I_{1}$.
@@
Inductively, one constructs $I_{1}\supset I_{2}\supset\ldots$ a decreasing sequence of sub-chainable compact set such that any point $x\in I_{n}$ satisfies that from time 0 to time $t_{1}+\ldots+t_{n}$ the $G_{t}$-orbit of $x$ does not intersect $D^{\prime}_{i}\cup\ldots\cup D^{\prime}_{N}$. It follows that the positive orbit of every $x$ in $\bigcap_{n}I_{n}$ does not intersect $D^{\prime}_{i}\cup\ldots\cup D^{\prime}_{N}$. Then it suffices to take $K$ equal to the omega limit of a point in $\bigcap_{n}I_{n}$ to obtain a compact $G_{t}$-invariant subset of $\Lambda_{f}$ disjoint from $\{x_{1},\ldots,x_{N}\}$. This ends the proof of the claim. ∎
@@
To finish the proof of Theorem 2.8 we can now argue as follows. Let $\Lambda_{1}\subset\Lambda_{f}$ be a $G_{t}$-minimal set and $x_{1}$ be a point in $\Lambda_{1}$. By Claim 3.11 there exists $K_{2}\subset\Lambda_{f}$ a compact $G_{t}$-invariant set such that $K_{2}\cap\{x_{1}\}=\emptyset$. Let $\Lambda_{2}\subset K_{2}$ be a $G_{t}$-minimal set. Note that $\Lambda_{1}$ and $\Lambda_{2}$ are disjoint (otherwise the point $x_{1}$ would lie in the intersection $\Lambda_{1}\cap\Lambda_{2}$, which is impossible).

### v2

@@ Abstract.
Keywords: Quasigeodesic flows, 3-manifolds, foliations, rotation theory, minimal homeomorphisms.
MSC 2022: Primary: 57R30, 37C27; Secondary: 53C12, 53C23.
@@ 1. Introduction
Let $f:M\to M$ be a minimal homeomorphism of a closed hyperbolic 3-manifold homotopic to the identity. Then, there is a $G_{\delta}$-dense [^1] I.e. containing a countable intersection of open and dense sets. set of points $\mathcal{G}\subset M$ so that if $x\in\mathcal{G}$ and $\tilde{x}\in\widetilde{M}$ is a lift of $x$ then:
@@
For flows, in [Fra] more information than that given by Theorem A is obtained about these compact sets (one can show these are periodic orbits of the flow), but such a result cannot hold for homeomorphisms (see Example 5.27). One can still expect to say more about the structure of these sets, but we achieve this only for some particular class of quasi-geodesic homeomorphisms. The following result is the technical core of this paper and extends ideas going back to [Ha] which were developed in some particular cases in [BFFP]. See also the recent [FrLa] for further important developments.
@@
Here we need to precise some terms. Given a foliation $\mathcal{F}$ of a closed 3-manifold, we say it is uniform-$\mathbb{R}$-covered if when lifted to the universal cover, we have that $\widetilde{\mathcal{F}}$ verifies that every leaf is a properly embedded plane in $\widetilde{M}$ and for every pair of leaves $L,L^{\prime}$ we have that the Hausdorff distance between $L$ and $L^{\prime}$ is bounded (see [FP1] for discussion, and below we give some more equivalences of this notion). Such foliations are quite abundant in hyperbolic 3-manifolds and extend the notion of fibrations for fibered hyperbolic 3-manifolds: they are always blow ups of slitherings (see [Thu2, Cal2]). The reason for calling them $\mathbb{R}$-covered is because in particular, we can see that the leaf space $\mathcal{L}_{\mathcal{F}}=\widetilde{M}$$_{\mathcal{F}}$ is homeomorphic to $\mathbb{R}$.
We say that $f$ has positive escape rate with respect to $\mathcal{F}$ if for every $x\in\widetilde{M}$ and a good lift $\widetilde{f}$ of $f$ we have that the leaf of $\widetilde{f}^{n}(x)$ goes to $\infty$ in $\mathcal{L}_{\mathcal{F}}$.
As an application of Theorem C, in § 6 we show that if a homeomorphism has positive speed with respect to some homological direction then it must have many invariant sets. We note here that the invariant sets that we produce have all at least topological dimension one [^2] This is optimal, one cannot ensure that the invariant sets are smaller, or that there are periodic orbits. For instance, one can consider the suspension flow of a pseudo-Anosov homeomorphism of a surface and take $f$ to be an irrational time of this suspension. The smallest closed invariant sets are circles, and the homeomorphism is quasi-geodesic (see Proposition 2.12). (see Proposition 5.26). In Example 5.27 we show that these invariant sets can still be quite wild.
@@
Acknowledgements: Part of this paper is the content of the master thesis [Go] of the first author, made in PEDECIBA-Udelar in Uruguay. The authors would like to thank Ian Agol, Alfonso Artigue, Jairo Bochi, Sylvain Crovisier, Sergio Fenley, Pablo Lessa, Ana Rechtman, Jonathan Zung for useful comments and exchange. We thank the referee for some insightful comments, including some questions that we have included in the text.
@@ 2.1. Escape rate
Consider the function $x\mapsto d(\widetilde{f}(x),x)$ defined on $\widetilde{M}$. Since $\widetilde{f}$ commutes with deck transformations, which act as isometries, this function is $\pi_{1}(M)$ equivariant (and thus defines a function $\varphi:M\to\mathbb{R}_{\geq 0}$). Similarly, we can define $x\mapsto d(\widetilde{f}^{n}(x),x)$ which induces a function $\varphi^{(n)}:M\to\mathbb{R}_{\geq 0}$. The sequence $\varphi^{(n)}$ is subaditive, that is $\varphi^{(n+m)}\leq\varphi^{(n)}\circ f^{m}+\varphi^{(m)}$ by the triangle inequality. This implies that for every ergodic $f$-invariant measure $\mu$ we have a well defined escape rate defined as $\ell_{\mu}=\inf_{n}\frac{1}{n}\int\varphi^{(n)}d\mu$, which coincides with the limit $\ell_{\mu}=\lim_{n}\frac{1}{n}\varphi^{(n)}(x)$ for $\mu$-almost every $x\in M$ (see [GGL, §1.2] for a detailed presentation, we will use the facts presented there below).
@@ 2.2. Homological rotation
The proof of this theorem is an application of Theorem C together with some properties of surfaces in hyperbolic 3-manifolds. As noted by the referee, it could be that the previous result holds only assuming that for every ergodic $\mu$ one has that $\ell_{c}(\mu)\neq 0$ (it certainly does if one assumes this for every invariant measure thanks to ergodic decomposition), but we have not been able to show this nor to produce a counterexample. One should also point out that in the case of flows, thanks to [Sh] the assumptions imply that the flow is a suspension of some map in a surface, and since $M$ is hyperbolic the map is homotopic to pseudo-Anosov and this is enough to conclude. Here, the challenge is to work with homeomorphisms.
@@
[…]
@@
[…]
@@
If $f:M\to M$ is a minimal homeomorphism of a closed $3$-manifold homotopic to the identity and $\widetilde{f}:\widetilde{M}\to\widetilde{M}$ is a good lift, then, for every $x\in\widetilde{M}$ we have that the sequence $d(\widetilde{f}^{n}(x),x)$ is unbounded. In fact, if for some point $x$ this sequence is bounded, we can consider the closure of its $\widetilde{f}-$orbit, the boundary of which is a compact $\widetilde{f}-$invariant set with empty interior. Thus, its projection to $M$ is a closed, proper, $f$-invariant set, contradicting minimality. In the known examples in hyperbolic 3-manifolds, minimal homeomorphisms verify that every point has escape rate of the order of $d(\widetilde{f}^{n}(x),x)\sim\log n$. We do not know if it is possible to expect that every minimal homeomorphism of a hyperbolic 3-manifold has escape rate slower than $n^{\alpha}$ for some (every) $\alpha>0$.
@@ 2.3. The quasi-geodesic case
This will be proved in § 3. Let us mention that Theorem 2.8 (as well as Theorem A) are true in higher dimensions with the same proof.
@@ 2.4. $\mathbb{R}$-covered foliations
$B(Y,\varepsilon)$ $=\{x\in\widetilde{M}:d(x,Y)<\varepsilon\}$
$B_{L}(X,\varepsilon)$ $=\{x\in L:d_{L}(x,X)<\varepsilon\}.$
If the foliation has no torus leaves (more generally, if it is Reebless) we know that the leaf space $\mathcal{L}=\widetilde{M}$$_{\widetilde{\mathcal{F}}}$ is a simply connected (possibly non-Hausdorff) one-dimensional manifold. We say that $\mathcal{F}$ is $\mathbb{R}$-covered if $\mathcal{L}$ is Hausdorff (equivalently, homeomorphic to $\mathbb{R}$). We say that $\mathcal{F}$ is uniform if for every pair of leaves $L,L^{\prime}\in\widetilde{\mathcal{F}}$ there is $C>0$ such that $L$ is contained in the $C$-neighborhood of $L^{\prime}$ and viceversa. See [Cal2, Chapter 9] for more on these foliations, in particular we will use [Cal2, Thm. 9.15 and Lemma 9.10].
@@
If $\widetilde{\mathcal{F}}$ is $\mathbb{R}$-covered and uniform, there exists a homeomorphism $Z:\mathcal{L}\to\mathcal{L}$ that commutes with the action of $\pi_{1}(M)$ on $\mathcal{L}=\widetilde{M}$$\widetilde{\mathcal{F}}$, and a constant $c>0$ such that for every leaf $L\in\widetilde{\mathcal{F}}$, we have the bound $d(L,Z(L))>c$.
@@ 3. Quasi-geodesic homeomorphisms
In this section we consider a homeomorphism $f:M\to M$ of a hyperbolic 3-manifold which is homotopic to the identity and such that it is quasi-geodesic in the sense of §2.3. The goal of this section is to prove Theorem 2.8 and that such homeomorphisms must have positive topological entropy. Together, this will imply Theorem A. We note here that while all this section is stated in the 3-dimensional case, it works equally well for homeomorphisms of hyperbolic $n$-manifolds (or even manifolds admitting negatively curved metrics), we leave the verification of this to the interested reader.
@@ 3.1. Constructing the invariant set
Note that we do not claim that the image of these maps is closed, so they need not be surjective. These properties are analogous to some properties verified for quasi-geodesic flows shown in [Cal2, §10.7]. A more detailed study of the maps $e^{\pm}$ when one studies a quasi-geodesic flow rather than a homeomorphism can be found in [Fra, FrLa].
@@
While the association $f\mapsto\Lambda_{f}$ may not be locally constant (even in the case of quasi-geodesic flows), it could be that, as in the case of flows, there is a geometric core which is independent of the map as long as one varies $f$ continuously (see [FrLa] for precise statements in the case of flows). It could be interesting to study this further for homeomorphisms. For our purposes, we will only need the following:
@@
This is a direct consequence of the properties we proved for the functions $e^{+}$ and $e^{-}$ in Lemmas 3.1 and $3.$ ∎
@@ 3.2. Invariant subsets
Since $\gamma_{n}\to\infty$ in $\pi_{1}(M)$ (this is equivalent to saying that the norm $\gamma_{n}\in\mathrm{Isom}(\mathbb{H}^{3})\cong\mathrm{PSL}_{2}(\mathbb{C})$ goes to infinity) we know that there is a sequence of neighborhoods $U_{n}\subset\partial\mathbb{H}^{3}$ shrinking to $v_{+}$ (with the induced metric of $\mathbb{R}^{3}$) whose complement is mapped in some open set $V_{n}$ of diameter going to 0 (again, with the induced metric of $\mathbb{R}^{3}$). We can choose also $U_{n}$ and $V_{n}$ so that the action of $\gamma_{n}$ expands all vectors tangent to $U_{n}$. Also, since $v_{+}\neq v_{-}$, for large enough $n$ we know that $v_{-}\notin U_{n}$. By cutting $\eta_{0}$ if necessary, we can assume that $e^{-}(\eta_{0}([0,1])\cap U_{n}=\emptyset$. Up to taking a subsequence, we know that $V_{n}\to\xi\in\partial\mathbb{H}^{3}$ and thus we get that $\gamma_{n}e^{-}(\eta_{0}([0,1]))=e^{-}(\gamma_{n}\eta_{0}([0,1]))\to\xi$.
@@
Let $D_{i}:=\mathcal{W}^{u}_{10\epsilon}(\mathcal{W}^{s}_{10\epsilon}(x_{i}))$ for every $i\in\{1,\ldots,N\}$, for some $\epsilon\in(0,\epsilon_{0})$ small enough so that if $i\neq j$, then $D_{i}\cap D_{j}=\emptyset$ and, moreover, if a point $x\in D_{i}$ satisfies that $G_{t}(x)\in D_{j}$ for some $t>0$, then $t\geq T$. Moreover, let $\epsilon\in(0,\epsilon_{0})$ be small enough so that $I$ admits a sub-chainable continua $J_{1}$ of diameter $\epsilon$. Also let $D^{\prime}_{i}:=\mathcal{W}^{u}_{\epsilon/2}(\mathcal{W}^{s}_{\epsilon/2}(x_{i}))$ for every $i\in\{1,\ldots,N\}$.
@@
Note that $\operatorname{diam}(G_{t}(J_{1}))$ tends to infinity with $t$ since $J_{1}$ is contained in a strong unstable manifold of $G_{t}$. Let $t_{1}>0$ be the first positive time such that one of the following two conditions happen: either $\operatorname{diam}(G_{t_{1}}(J_{1}))=3\epsilon$ or $G_{t_{1}}(J_{1})\cap D^{\prime}_{i_{1}}\neq\emptyset$ for some ${i_{1}}\in\{1,\ldots,N\}$. If the former happens, let $J_{2}\subset G_{t_{1}}(J_{1})$ be a sub-chainable continua such that $\operatorname{diam}(J_{2})=\epsilon$. If the latter happens (i.e. $G_{t_{1}}(J_{1})\cap D^{\prime}_{i_{1}}$ while $\mathrm{diam}(G_{t_{1}}(J_{1}))\leq 3\varepsilon$), note that $G_{t_{1}}(J_{1})\subset D_{i_{1}}$, and let $J_{2}\subset G_{t_{1}}(J_{1})$ be a sub-chainable continua disjoint from $D_{i_{1}}^{\prime}$ such that $\operatorname{diam}(J_{2})=\epsilon$. Let $I_{1}=J_{1}$ and $I_{2}:=G_{-t_{1}}(J_{2})$. Note that $I_{2}\subset I_{1}$ and that for $x\in I_{2}$ one has that $G_{t}(x)\notin D^{\prime}_{1}\cup\ldots\cup D_{N}^{\prime}$ for $0\leq t\leq t_{1}$.
@@
Now, let $t_{2}>0$ be the first positive time such that, again, one of the following two conditions happen: either $\operatorname{diam}(G_{t_{2}}(J_{2}))=3\epsilon$ or $G_{t_{2}}(J_{2})\cap D^{\prime}_{i_{2}}\neq\emptyset$ for some ${i_{2}}\in\{1,\ldots,N\}$. Note that $t_{2}\geq T$. If the former happens, let $J_{3}\subset G_{t_{2}}(J_{2})$ be a sub-chainable continua such that $\operatorname{diam}(J_{3})=\epsilon$. If the latter happens, note that $G_{t_{2}}(J_{2})\subset D_{i_{2}}$, and let $J_{3}\subset G_{t_{2}}(J_{2})$ be a sub-chainable continua disjoint from $D^{\prime}_{i_{2}}$ such that $\operatorname{diam}(J_{3})=\epsilon$ (note that in this case, this is possible since $\operatorname{diam}(G_{t_{2}}(J_{2}))\geq 2\epsilon$ is guaranteeed because of Claim 3.10 and the fact that $t_{2}\geq T$). Let $I_{3}:=G_{-t_{1}-t_{2}}(J_{3})$. Note that $I_{3}\subset I_{2}\subset I_{1}$ and that for $x\in I_{3}$ one has that $G_{t}(x)\notin D^{\prime}_{1}\cup\ldots\cup D_{N}^{\prime}$ for $0\leq t\leq t_{1}+t_{2}$..
@@
Inductively, one constructs $I_{1}\supset I_{2}\supset\ldots$ a decreasing sequence of non empty sub-chainable compact set such that any point $x\in I_{n}$ satisfies that from time 0 to time $t_{1}+\ldots+t_{n}\geq(n-1)T$ the $G_{t}$-orbit of $x$ does not intersect $D^{\prime}_{1}\cup\ldots\cup D^{\prime}_{N}$. Since $I_{n}$ are nested non empty compact sets, one knows that $\bigcap_{n}I_{n}\neq\emptyset$. It follows that the positive orbit of every $x\in\bigcap_{n}I_{n}$ does not intersect $D^{\prime}_{1}\cup\ldots\cup D^{\prime}_{N}$. Then it suffices to take $K$ equal to the omega limit of a point in $\bigcap_{n}I_{n}$ to obtain a compact $G_{t}$-invariant subset of $\Lambda_{f}$ disjoint from $\{x_{1},\ldots,x_{N}\}$. This ends the proof of the claim. ∎
@@
To finish the proof of Theorem 2.8 we can now argue as follows. Let $\Lambda_{1}\subset\Lambda_{f}$ be a $G_{t}$-minimal set and $x_{1}$ be a point in $\Lambda_{1}$. By Claim 3.11 there exists $K_{2}\subset\Lambda_{f}$ a compact $G_{t}$-invariant set such that $K_{2}\cap\{x_{1}\}=\emptyset$. Let $\Lambda_{2}\subset K_{2}$ be a $G_{t}$-minimal set. Note that $\Lambda_{1}$ and $\Lambda_{2}$ compact minimal sets, thus, they must be disjoint (becase minimal sets are disjoint or equal and $x_{1}\in\Lambda_{1}\setminus\Lambda_{2}$).